OpenStudy (anonymous):

m=4/5, (3, -4) answer needs to be in slope-intercept form

OpenStudy (anonymous):

Do you know how you would put it in point-slope form?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

(y +4) = 4/5 (x-3) => 5y + 20 = 4x - 12 => 5y = 4x -32 => y = 4/5 x - 32/5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

u welcome...... use it also : www.algebra.com